OpenStudy (anonymous):

Indices question; simplify: (64)^-(5/6)

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

what comes to your mind first?

OpenStudy (anonymous):

so because its a negative it will be 1 over the positive

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

i now get that bit i dont know how to figure the answer..

OpenStudy (anonymous):

you have to multiply the powers

OpenStudy (triciaal):

how can you rewrite 64 as a number raised to a power

OpenStudy (anonymous):

would it be 2^6

OpenStudy (triciaal):

ask what are the prime factors of 64 do you see that 64 = 2^6

OpenStudy (triciaal):

great!

OpenStudy (anonymous):

Yes!! :))

OpenStudy (triciaal):

so what you have 64^(-5/6) = 2^6^(-5/6) you have a power raised to a power so what do you do ?

OpenStudy (anonymous):

multiply them

OpenStudy (triciaal):

yes go ahead

OpenStudy (anonymous):

would you do 64/2

OpenStudy (anonymous):

32

OpenStudy (triciaal):